# A strike of luck: could the KM3–230213A event be caused by an evaporating primordial black hole?

### v1

@@ Abstract
We investigate whether the ultra high energy neutrino inferred by the recent KM3NeT observation could have originated from an evaporating black hole. Given the characteristics of black hole (BH) evaporation mechanism, any object capable of producing particles in the energy range of the detected event ($\sim$100-800 PeV) must have a mass $M_{\rm BH}\lesssim 10^{7}\rm g$. No known astrophysical mechanism can generate black holes of such low mass, leaving primordial black holes (PBHs) –potentially formed at the end of cosmic inflation– as the only viable candidates. Black holes with masses $M_{\rm BH}\lesssim 10^{7}\rm g$ have lifetimes shorter than $10^{-5}$ seconds, meaning PBHs in this mass range should have fully evaporated by now. However, recent studies suggest that quantum effects, collectively referred to as the “memory burden”, may slow down black hole evaporation, potentially extending the lifetimes of low–mass PBHs to timescales comparable to or exceeding the Hubble time. We systematically explore the parameter space of memory-burdened PBHs, assuming that they constitute a fraction of the dark matter ($f_{\rm PBH}$) within current constraints, and identify viable regions that could explain the KM3–230213A event. We further predict the occurrence rate of similar events, and find that KM3NeT’s current configuration could test this scenario within a few years.
@@
The standard semi-classical picture of black hole evaporation assumes that the object remains classical throughout its lifetime Hawking (1975). However, growing evidence suggests that this framework may be incomplete, particularly in addressing the information loss paradox Preskill (1992). A key limitation of Hawking’s derivation is the neglect of back-reaction effects—how the emission of radiation influences the quantum state of the black hole itself. This omission becomes critical when the energy of emitted particles approaches the total energy of the black hole.
Recent research Dvali (2018); Dvali et al. (2020, 2024) suggests that such back-reaction effects give rise to a universal “memory burden”. This phenomenon emerges because the information encoded in a system resists its decay, driven by the response of quantum modes tied to its entropic degrees of freedom. As a result, once a black hole’s mass drops below a certain threshold, back-reaction significantly alters its evolution, slowing the evaporation process and extending its lifetime. This opens the possibility that black holes with masses $M_{\rm PBH}\lesssim 10^{15}$ g may still be evaporating today, with potential observational consequences Balaji et al. (2024); Barman et al. (2024a); Bhaumik et al. (2024); Barman et al. (2024b); Kohri et al. (2024); Jiang et al. (2024).
@@
Due to the mechanism described, their mass function is almost unconstrained “a priori”, and over the last decade this has caused consistent activity about the possibility that they could constitute all or a fraction of the Dark Matter, Cole et al. (2023) and references therein. Constraints on this occurrence arise somewhat naturally for those PBHs which would have had to evaporate entirely by the Hubble time Boccia et al. (2024); Luo et al. (2021); Kawasaki et al. (1999, 2000, 2005, 2018, 2020); Hasegawa et al. (2019); de Salas et al. (2015), namely those whose masses don’t exceed $M_{\rm PBH}\gtrsim 10^{15}\rm g$, whereas other types of physical mechanisms strongly constrain those masses that should survive the Universe’s lifetime Carr et al. (2021).
The potential “life-extending” effect of memory burden with respect to a standard scenario has prompted a reassessment of the potential observations and of the constraints derived. It has also opened up to the enticing possibility that extremely high energy particles may be injected in the local Universe by the evaporation of PBHs of low masses, suppressed in the standard case Thoss et al. (2024); Chianese et al. (2025); Zantedeschi and Visinelli (2024).
@@
$M_{\mathrm{PBH}}\,[\mathrm{g}]$ $k$ $f_{\mathrm{PBH}}$ $t^{1\nu}_{\rm exp}$ [yr]
$2.5\times 10^{4}$ 1.8 $5.6\times 10^{-8}$ $4.3$
$1.0\times 10^{5}$ 1.6 $3.2\times 10^{-8}$ $5.0$
$4.0\times 10^{5}$ 1.4 $8.8\times 10^{-9}$ $4.2$
$5.0\times 10^{5}$ 1.4 $3.5\times 10^{-8}$ $3.9$
$6.3\times 10^{5}$ 1.4 $1.4\times 10^{-7}$ $3.7$
$2.5\times 10^{6}$ 1.2 $9.1\times 10^{-9}$ $4.3$
$3.1\times 10^{6}$ 1.2 $2.6\times 10^{-8}$ $5.6$
$4.0\times 10^{6}$ 1.2 $6.7\times 10^{-8}$ $7.7$
$2.0\times 10^{7}$ 1.0 $6.8\times 10^{-9}$ $20$
$2.5\times 10^{7}$ 1.0 $3.5\times 10^{-8}$ $21$
Table 1: Exposure time for one neutrino event $(t^{1\nu}_{\rm exp})$ to be expected within the current KM3NeT configuration for the ten most highest-yielding pairs of $(M_{\mathrm{PBH}},k)$, assuming they constitute the highest dark matter fraction $f_{\mathrm{PBH}}$ allowed by current constraints.
Discussion and forecasts In light of the caveats introduced, we note at this point that existing constraints limit $f_{PBH}\ll 1$ for all the parameter combinations found. Namely: none of the pairs of $M_{\mathrm{PBH}}$ and $k$ found to produce a sizable flux can form the entirety of the dark matter, and the number counts shown in the plot must be appropriately rescaled for each point in Figure 1. This leads –for all combinations identified, for the actual exposure time of the event observed (335 days), and for the directional characteristics of the event– to an expected event number smaller than one.
@@
We assume that the KM3–230213A event has indeed been produced by a neutrino emitted by an evaporating black hole of low mass, survived until today owing to the memory-burden effect. Within this scenario, it is now possible to make predictions of the number of future events expected within the KMN3NeT detector in the future, for each mass $M_{\mathrm{PBH}}$ and $k$ identified, and assuming that their contribution to the dark matter fraction is maximal, i. e.  the highest $f_{\rm PBH}$ allowed by the current constraints. The latter, presented in Table 1, are derived from the results of Thoss et al. (2024); Chianese et al. (2025). They were established by analyzing the absence of detected $\gamma$-rays, and neutrinos within the IceCube detector, which would otherwise be expected from the evaporation of primordial black holes.
@@
We apply the methodology described above to perform this analysis, presenting the results in Table 1. We show there the ten combinations ($M_{\mathrm{PBH}}$, $k$) with the highest expected event counts, assuming that each specific pair of black hole mass $M_{\mathrm{PBH}}$ and memory burden strength $k$ contributes the maximum permissible fraction of Dark Matter $f_{\mathrm{PBH}}$ within existing constraints, and report the corresponding exposure time required to expect one neutrino event within the KM3NeT. It is to be noted here that the configuration of the experiment is constantly changing as new detectors are added, with the effect of enhancing the effective area with a rate that is not possible to predict reliably. Our estimates are therefore to be regarded as conservative, and to be updated as the experiment keeps being upgraded. Given these uncertainties, we refrain from presenting more detailed detection forecasts –such as the angular dependence of the signal due to the orientation of the densest dark matter regions in the Galaxy, or exclusion plots based on predicted sensitivities– as they fall beyond the scope of this study. A more comprehensive study should be undertaken once the updated instrument configurations become available.
@@
Conclusions Within a cosmological framework where primordial black holes (PBHs) constitute a fraction of dark matter ($f_{\rm PBH}$), we have investigated whether PBHs evaporating today could have produced the KM3–230213A event by emitting a spectrum of particles, including neutrinos. For this scenario to be viable, their lifetime must be extended by the recently proposed memory-burden effect. We have identified the PBH masses $M_{\mathrm{PBH}}$ and the memory burden effect strength parameter $k$ that maximize the number of expected events in the KM3NeT detector while remaining consistent with existing observational constraints. Although the probability that KM3–-230213A was produced by an evaporating memory-burdened PBH is low ($\ll 1$) for all identified parameter sets ($M_{\mathrm{PBH}}$, $k$) identified, one can nevertheless take an optimistic stance and put this hypothesis to the test.
@@
Following this approach, we predict the number of neutrino events expected in KM3NeT under the assumption that PBHs contribute the maximum fraction of dark matter $f_{\rm PBH}$ allowed by current constraints. The predicted event rate is significant, reaching the order of unity within a few years. This would make it possible to either confirm or rule out this scenario, subjecting the memory-burdened PBH paradigm to imminent scrutiny with neutrino observations.

### v2

@@ Abstract
We investigate whether the ultra high energy neutrino inferred by the recent KM3NeT observation could have originated from an evaporating black hole. Given the characteristics of black hole (BH) evaporation mechanism, any object capable of producing particles in the energy range of the detected event ($\sim$100-800 PeV) must have a mass $M_{\rm BH}\lesssim 10^{7}\rm g$. No known astrophysical mechanism can generate black holes of such low mass, leaving primordial black holes (PBHs) –potentially formed at the end of cosmic inflation– as the only viable candidates. Black holes with masses $M_{\rm BH}\lesssim 10^{7}\rm g$ have lifetimes shorter than $10^{-5}$ seconds, meaning PBHs in this mass range should have fully evaporated by now. However, recent studies suggest that quantum effects, collectively referred to as the “memory burden”, may slow down black hole evaporation, potentially extending the lifetimes of low–mass PBHs to timescales comparable to or exceeding the Hubble time. We systematically explore the parameter space of memory-burdened PBHs, assuming that they constitute a fraction of the dark matter ($f_{\rm PBH}$) within current constraints, and identify viable regions that could explain the KM3–230213A event. We further predict the occurrence rate of similar events, and find that KM3NeT and future neutrino experiments could test this scenario over the next years.
@@
The standard semi-classical picture of black hole evaporation assumes that the object remains classical throughout its lifetime Hawking (1975). However, growing evidence suggests that this framework may be incomplete, particularly in addressing the information loss paradox Preskill (1992). A key limitation of Hawking’s derivation is the neglect of back-reaction effects, how the emission of radiation influences the quantum state of the black hole itself. This omission becomes critical when the energy of emitted particles approaches the total energy of the black hole.
Recent research Dvali (2018); Dvali et al. (2020, 2024); Davies et al. (2024) suggests that such back-reaction effects give rise to a universal “memory burden”. This phenomenon emerges because the information encoded in a system resists its decay, driven by the response of quantum modes tied to its entropic degrees of freedom. As a result, once a black hole’s mass drops below a certain threshold, back-reaction significantly alters its evolution, slowing the evaporation process and extending its lifetime. This opens the possibility that black holes with masses $M_{\rm PBH}\lesssim 10^{15}$ g may still be evaporating today, with potential observational consequences Balaji et al. (2024); Barman et al. (2024a); Bhaumik et al. (2024); Barman et al. (2024b); Kohri et al. (2024); Jiang et al. (2024).
@@
Due to the mechanism described, their mass function is almost unconstrained “a priori”, and over the last decade this has caused consistent activity about the possibility that they could constitute all or a fraction of the Dark Matter, Cole et al. (2023) and references therein. Constraints on this occurrence arise somewhat naturally for those PBHs which would have had to evaporate entirely by the Hubble time Boccia et al. (2024); Luo et al. (2021); Kawasaki et al. (1999, 2000, 2005, 2018, 2020); Hasegawa et al. (2019); de Salas et al. (2015); Wu and Xu (2024), namely those whose masses don’t exceed $M_{\rm PBH}\gtrsim 10^{15}\rm g$, whereas other types of physical mechanisms strongly constrain those masses that should survive the Universe’s lifetime Carr et al. (2021).
The potential “life-extending” effect of memory burden with respect to a standard scenario has prompted a reassessment of the potential observations and of the constraints derived. It has also opened up to the enticing possibility that extremely high energy particles may be injected in the local Universe by the evaporation of PBHs of low masses, suppressed in the standard case Thoss et al. (2024); Chianese et al. (2025); Zantedeschi and Visinelli (2024); Loc (2025).
@@
$M_{\mathrm{PBH}}\,[\mathrm{g}]$ $k$ $f_{\mathrm{PBH}}$ $t_{\mathrm{KM}}\,[\mathrm{yr}]$ $t_{\mathrm{IC2}}\,[\mathrm{yr}]$
$2.0\times 10^{6}$ 1.4 $2.1\times 10^{-5}$ $21$ $5.2\times 10^{-1}$
$2.0\times 10^{6}$ 1.6 $8.6\times 10^{-1}$ $21$ $5.2\times 10^{-1}$
$4.0\times 10^{6}$ 1.2 $2.8\times 10^{-8}$ $19$ $3.7\times 10^{-1}$
$4.0\times 10^{6}$ 1.4 $1.5\times 10^{-3}$ $19$ $3.7\times 10^{-1}$
$5.0\times 10^{6}$ 1.2 $1.6\times 10^{-7}$ $12$ $2.1\times 10^{-1}$
$5.0\times 10^{6}$ 1.4 $8.7\times 10^{-3}$ $13$ $2.2\times 10^{-1}$
$6.3\times 10^{6}$ 1.2 $3.6\times 10^{-7}$ $20$ $2.9\times 10^{-1}$
Table 1: Exposure time for one neutrino event $t_{\mathrm{KM}}$ to be expected within the current KM3NeT configuration for the most highest-yielding pairs of $(M_{\mathrm{PBH}},k)$, assuming they constitute the highest dark matter fraction $f_{\mathrm{PBH}}$ allowed by current constraints. We also report forecast times for one neutrino event to be expected within the forthcoming IceCube Gen-2 detector, $t_{\mathrm{IC2}}$.
Discussion and forecasts In light of the caveats introduced, we note at this point that existing constraints limit $f_{\rm PBH}\ll 1$ for all the parameter combinations found. Namely: none of the pairs of $M_{\mathrm{PBH}}$ and $k$ found to produce a sizable flux can form the entirety of the dark matter, and the number counts shown in the plot must be appropriately rescaled for each point in Figure 1. This leads –for all combinations identified, for the actual exposure time of the event observed (335 days), and for the directional characteristics of the event– to an expected event number smaller than one.
@@
We assume that the KM3–230213A event has indeed been produced by a neutrino emitted by an evaporating black hole of low mass, survived until today owing to the memory-burden effect. Within this scenario, it is now possible to make predictions of the number of future events expected within the KM3NeT detector in the future, for each mass $M_{\mathrm{PBH}}$ and $k$ identified, and assuming that their contribution to the dark matter fraction is maximal, i. e.  the highest $f_{\rm PBH}$ allowed by the current constraints. The latter have been obtained by the absence of neutrino events with the correct characteristics within the IceCube detector in a seven year dataset Chianese et al. (2025), and that of $\gamma$-rays within several $\gamma$ observatories Thoss et al. (2024); Chianese (2025). By adopting the values of $f_{\rm PBH}$ shown in Table 1 into Eq.12, we are ensuring that our working hypothesis does not violate the lack of previous observations. It is worth noting here that the tension between the KM3NeT and the IceCube data found by Li et al. (2025) does not apply at face value to our scenario, as it has been derived for cosmogenic neutrinos and an isotropic signal. This is not the case for the PBH–induced neutrino flux, which follows a Galactic DM distribution.
@@
We apply the methodology described above to perform this analysis, presenting the results in Table 1. We show there the combinations ($M_{\mathrm{PBH}}$, $k$) with the highest expected event counts, assuming that each specific pair of black hole mass $M_{\mathrm{PBH}}$ and memory burden strength $k$ contributes the maximum permissible fraction of Dark Matter $f_{\mathrm{PBH}}$ within existing constraints, and report the corresponding exposure time required to expect one neutrino event within the KM3NeT. It is to be noted here that the configuration of the experiment is constantly changing as new detectors are added, with the effect of enhancing the effective area with a rate that is not possible to predict reliably. Our estimates are therefore to be regarded as conservative, and to be updated as the experiment keeps being upgraded. Given these uncertainties, we refrain from presenting more detailed detection forecasts –such as the angular dependence of the signal due to the orientation of the densest dark matter regions in the Galaxy, or exclusion plots based on predicted sensitivities– as they fall beyond the scope of this study. A more comprehensive study should be undertaken once the updated instrument configurations become available. We also report forecasts for the exposure time in the IceCube Gen-2 detector –currently under construction, and with effective area estimated as from Aartsen et al. (2021)– to show the potential of future generation detectors. Also, gamma-ray observations should accompany future neutrino ones, so specific forecasts should be performed in order to test, or constrain, the hypothesis put forward in this paper via an appropriate multi-messenger strategy for the entire memory burdened PBH parameter space.
@@
Conclusions Within a cosmological framework where primordial black holes (PBHs) constitute a fraction of dark matter ($f_{\rm PBH}$), we have investigated whether PBHs evaporating today could have produced the KM3–230213A event by emitting a spectrum of particles, including neutrinos. For this scenario to be viable, their lifetime must be extended by the recently proposed memory-burden effect. We have identified the PBH masses $M_{\mathrm{PBH}}$ and the memory burden effect strength parameter $k$ that maximize the number of expected events in the KM3NeT detector while remaining consistent with existing observational constraints. Although the probability that KM3–230213A was produced by an evaporating memory-burdened PBH is very low ($\ll 1$) for all identified parameter sets ($M_{\mathrm{PBH}}$, $k$) identified, one can nevertheless take an optimistic stance and put this hypothesis to the test.
@@
Following this approach, we predict the number of neutrino events expected in KM3NeT and Icecube Gen-2 under the assumption that PBHs contribute the maximum fraction of dark matter $f_{\rm PBH}$ allowed by current constraints. The predicted event rate is significant, reaching the order of unity within a few years. This would make it possible to either confirm or rule out this scenario, subjecting the memory-burdened PBH paradigm to imminent scrutiny with neutrino observations.